OpenStudy (anonymous):

If ax*x + bx +c = 0, then what is x?

ganeshie8 (ganeshie8):

heard of quadratic formula before ?

OpenStudy (anonymous):

(a)x*x + (b)x + (c) = 0 d*d = b*b - 4ac (-b +/- d)/2a=x

OpenStudy (anonymous):

did i do that right?

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

thanks for your help :D

ganeshie8 (ganeshie8):

np :) generally we call d = b^2-4ac

ganeshie8 (ganeshie8):

so, \(\large x = \frac{-b\pm \sqrt{d}}{2a}\)

OpenStudy (anonymous):

OHHH i see now :D

ganeshie8 (ganeshie8):

its okay to call d^2 = b^2-4ac, if u feel thats more neat :)

OpenStudy (anonymous):

alright :D thanks again

ganeshie8 (ganeshie8):

yu wlcme :)